# Emotional Strain and Frustration Drivers in LLM-Human Interactions in Software Engineering

### v1

@@
Abstract
Large Language Models (LLMs) are increasingly integrated into various daily tasks in Software Engineering. Even though LLMs are widely used, some interactions can lead to unexpected challenges, which can cause emotions that can evolve into frustration. Frustration can negatively impact the engineer’s productivity and well-being if it escalates into stress and burnout, particularly given the demanding nature of Software Engineering. We explored software engineers’ emotional responses, specifically strains, and identified common causes of frustration when interacting with LLMs at work. We collected 62 survey responses from software practitioners and academics across various software companies and universities, using open-ended, closed-ended, and Likert-scale questions. Then, we analysed the survey responses using content analysis and descriptive statistics.
Our findings revealed that a majority of software engineers experience frustrations or other emotions that can evolve into frustrations. More specifically, we found that the leading causes of frustrations aligned with the engineers’ expectations and were related mainly to the responses’ correctness and reliability. Moreover, we noticed that another major frustration trigger comes from uncritical problems, such as the LLM failing to meet personal preferences (e.g., the output structure) since it would require the engineer to refactor or make more modifications to the LLM output before using it.
This study not only explores frustration triggers when using LLMs but also proposes guidelines to improve the software engineers’ experience, aiming to minimise long-term consequences on mental health.
Index Terms:
@@ Abstract
I Introduction
Software Engineering is full of challenges, from fixing bugs to dealing with changing requirements. Recently, Large Language Models (LLMs) like ChatGPT and GitHub Copilot have been used by software engineers to assist them in performing various tasks including code generation, and quality assurance [asare2023github, lubos2024leveraging]. Current research focuses on understanding how practitioners aim to increase their productivity and make their work process more efficient by targeting LLMs to automate the generation of software artifacts or receive guidance on how to solve certain problems [khojah2024beyond, storey2016disrupting].
However, when certain challenges and limitations of LLMs hinder their effectiveness (e.g., not arriving at a helpful response), engineers may experience frustration [weisz2022better]. Frustrations cause by LLMs as technologies are a component of Techno-stress that impact the engineers’ overall workflow, well-being, and productivity [tarafdar2007impact, muller2023can].
While frustrations have gotten little attention in LLM4SE research, we argue that understanding the causes of frustrations when using LLMs for software-related tasks is the first step to minimizing them and thus improving the productivity and well-being of practitioners both in the software engineering industry and academia.
In this study, we conduct an exploratory study to empirically investigate the causes of frustrations when software engineers interact with LLMs and propose ways to work towards improving productivity and well-being. We focus on the following research questions:
@@ I Introduction
Our study presents four main categories that can cause frustrating emotions for the software engineer. We found that the main cause of frustration is when the software engineers receive an unhelpful or incorrect answer, followed by misunderstanding the intention, failing to meet personal preferences and other limitations of LLMs. We argue that these categories are specific to software engineering since it is a domain that requires precise, context-aware, and technically accurate responses that can be directly applied, like using LLMs for code generation.
@@
We report that motivation can be temporarily reduced when certain expectations are unmet and frustration is caused. However, interestingly, we found that the software engineers’ motivation to continue with their tasks mostly remains unaffected by a frustrating interaction with LLMs. This was due to the urgency to complete the task (e.g., short sprints and deadlines), and the availability of alternatives (e.g., search engines or personal expertise).
RQ3: How can the experience be less frustrating?
We provide recommendations on areas to improve in LLMs that can be considered by chatbot designers. These recommendations align with the categories of different expectations of software engineers from LLMs. In addition, we argue for the need to provide training and raise awareness among software engineers (in academia and industry) in order to manage and minimize frustrations.
II Background and Related Work
@@
This sections presents the conception framework of this study, as well as, previous work done on the topic and the the research gap we are trying to address.
II-1 Large language models for Software Engineering
@@
With the rise of Large Language Models (LLMs), many industries have been undergoing a transition towards using them for various activities. In Software Engineering, recent research has shown that LLMs have the potential to support practitioners in a variety of tasks including, but not limited to implementation, testing, requirements engineering [dakhel2023github, wang2024software, norheim2024challenges, khojah2024beyond]. Moreover, despite the challenges LLMs impose on academia (e.g., bias and hallucination), they provide many opportunities for researchers and educators as assistants with creating study guides and academic writing [meyer2023chatgpt].
@@
Despite this extensive use and the time that LLMs can save when automating tasks, integrating LLMs into SE practices can lead to user frustration. Researchers have noted that while LLMs assist software practitioners, they may sometimes generate errors, causing them to spend additional time solving errors or seeking clarification. This can also trigger negative emotions, such as frustration [weisz2022better].
II-A Emotions
@@
In this study, we considered the definition by the American Psychological Association [apa_dictionary_emotion], “a complex reaction pattern, involving experiential, behavioural, and physiological elements, by which an individual attempts to deal with a personally significant matter or event”. We acknowledge the diversity of the conceptual definitions of emotions, we selected this definition for its scope, which integrates experiential, behavioural, and physiological dimensions. The Feeling Wheel, developed by Gloria Willcox [willcox1982feeling], is a psychological tool designed to facilitate identifying and expressing emotions. It is widely used in therapeutic settings and personal development to help individuals articulate their emotional states more precisely. The tool organises emotions into a concentric structure, with six core emotions—happy, sad, angry, scared, strong, and calm—situated at the center of the wheel, see the wheel in Figure1. As one moves outward from the core, these primary emotions subdivide into more specific and nuanced feelings, allowing for a more granular understanding of emotional experiences.
Figure 1: Willcox’s Emotions Wheel [neff2024feelings]
@@
In research contexts, the Feeling Wheel has been employed to categorise participant-reported emotions, enabling a systematic approach to the analysis of emotional expression. In this study, we used the adaptation done by [neff_emotional_2023] to explore how individuals experience their interaction with LLMs.
II-A1 Emotional Responses when Using Technology
@@
Salanova [salanova2014technostress] defines Technostrain as a negative psychological experience characterised by (1) heightened anxiety and fatigue (affective dimension), (2) scepticism (attitudinal dimension), and (3) a sense of inefficacy (cognitive dimension) associated with technology use. Aligned with this, Muller et al. [muller2023can] researched techno-frustration, a component of technostrain, which refers to the psychological strain caused by the disorganised or inefficient use of ICT. Techno-frustration describes the experience of feeling discouraged, uncertain, stressed, confused, and upset as a result of using ICT. Such psychological strain can lead to decreased job satisfaction or an increased risk of burnout [srivastava2015technostress]. Muller et al.’s work is the closest to our study; however, we focus on a specific interaction, LLM-user interaction, which is still underexplored despite the rapid integration of LLMs in software engineering.
@@
Regarding research on how technology impacts users’ emotions, studies have focused on the psychological effects of prolonged technology use, recognising that digital tools influence productivity and mental and emotional states (e.g.,[calvo2014positive, mark2015focused]). This focus reflects an effort to explore the balance between the benefits of technology and the potential strain it places on users, particularly in high-demand environments like software engineering. Examples of those efforts are in the form of interventions (e.g.[montes2024qualifying], [penzenstadler2022take]) or looking for the causes (e.g. [wong2023mental]).
@@
Furthermore, users’ expectations when using technology, particularly LLMs, are crucial to their frustration. Studies have shown that prior experiences with technology can shape users’ expectations, influencing their perceptions of performance and trust [ferreri2023identifying]. External factors, such as marketing and communication from technology developers, also play a critical role in setting these expectations [roscoe2017presentation]. When technology does not meet users’ expectations—whether based on past interactions or external portrayals—frustration can escalate, hindering effective task completion and satisfaction  [srivastava2015technostress]. This study aims to explore whether these established findings apply also to LLM interactions, and how this impacts software engineering tasks, a question currently underexplored in existing literature.
This study addresses a gap in the literature by exploring software engineers’ emotional responses, particularly their frustration when using large language models (LLMs), motivated by the growing use of LLMs in professional tasks.
This research aims to provide insights that can inform improvements in LLM design and user experience, ultimately enhancing software engineering productivity and reducing technostress.
III Methodology
The main goal of our study is to understand software engineers’ frustrating experiences when interacting with LLMs in order to be able to point out potential causes of frustrations and possible mitigation strategies. We are also interested in looking into other emotions than frustration that software engineers feel in different interactions and scenarios. We illustrate the different components of our methodology in Figure 2.
Figure 2: Exploratory Study Process.
III-A Study Design
This study employed an exploratory design to investigate the frustrations experienced by software engineers when using large language models (LLMs). Given the limited existing research on this specific topic, an exploratory approach was deemed appropriate to gain initial insights and identify patterns in how engineers interact with LLMs [stebbins2001exploratory]. Exploratory research is particularly useful when studying relatively new phenomena, as it allows for collecting qualitative and quantitative data without predefined hypotheses, facilitating a broad understanding of the issues at hand. We collected open and closed questions to ensure a comprehensive data set that addresses both specific aspects of LLM usage and participants’ broader, more subjective experiences.
III-B Data Collection
@@
In this study, we used a survey as our main data collection method. We initially designed the survey questions based on previous research about the usage of LLMs in Software Engineering and how it impacts their experience and productivity [khojah2024beyond] as well as general limitations of LLMs [hadi2023survey].
@@
After a pilot study with two PhD students (academia) and two software engineers (industry), we added more questions and refined the survey (available in our Zenodo package [zenodo]). In general, the survey includes (i) demographic questions and questions related to the participants’ experience and familiarity with LLMs, (ii) questions about their usage of LLMs during their work and their expectations from the LLMs, (iii) questions about the participants’ emotions when certain expectations are not met, (iv) open-ended questions about frustrations the participants experienced when using LLMs, and (v) Likert-scales of the level of importance of different LLMs abilities and aspects to minimize frustrating experiences.
@@
We created the survey on Google Forms and distributed it across social platforms (LinkedIn and Facebook). We also sent personal invitations to employees who work across 10 software organizations of different sizes (Startups and large companies) and domains (e.g., automotive and eLearning). We also invited researchers and academics in software engineering conferences (RE’24 and FSE’24) to allow our sample of software engineers to be diverse in terms of countries and domains.
III-C Data Analysis
@@
To analyse the open questions, we used content analysis with an inductive approach following the steps by Erlingsson and Brysiewicz [erlingsson2017hands]. The authors analysed the data systematically, aiming to generate patterns, codes and themes [lune2017qualitative]. We discussed each answer for each question, reaching a data agreement. The codes were then categorised based on similarities and differences [erlingsson2017hands]. For question eight, we used the emotion classification and feeling wheel by Willcox [willcox1982feeling] to categorise and identify the range of emotions expressed by participants, enabling a deeper understanding of their emotional responses during interactions with LLMs. For the Likert and close questions, we use descriptive statistics and create visualisations to improve readability. We could not perform a correlation analysis due to the size of our sample.
IV Results
IV-A Respondents Demographics

### v2

# Emotional Strain and Frustration in LLM Interactions in Software Engineering

@@
Abstract.
Large Language Models (LLMs) are increasingly integrated into various daily tasks in Software Engineering such as coding and requirement elicitation. Despite their various capabilities and constant use, some interactions can lead to unexpected challenges (e.g. hallucinations or verbose answers) and, in turn, cause emotions that develop into frustration. Frustration can negatively impact engineers’ productivity and well-being if they escalate into stress and burnout. In this paper, we assess the impact of LLM interactions on software engineers’ emotional responses, specifically strains, and identify common causes of frustration when interacting with LLMs at work. Based on 62 survey responses from software engineers in industry and academia across various companies and universities, we found that a majority of our respondents experience frustrations or other related emotions regardless of the nature of their work. Additionally, our results showed that frustration mainly stemmed from issues with correctness and less critical issues such as adaptability to context or specific format. While such issues may not cause frustration in general, artefacts that do not follow certain preferences, standards, or best practices can make the output unusable without extensive modification, causing frustration over time. In addition to the frustration triggers, our study offers guidelines to improve the software engineers’ experience, aiming to minimise long-term consequences on mental health.
@@ Abstract.
conference: The 29th International Conference on Evaluation and Assessment in Software Engineering; 17–20 June, 2025; Istanbul, Türkiyecopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Software Engineering (SE) comes with many challenges, from fixing bugs to dealing with changing requirements. Recently, Large Language Models (LLMs) and LLM-powered chatbots like ChatGPT and GitHub Copilot have been used by software engineers to assist them in performing various tasks including code generation, and quality assurance (Asare et al., 2023; Lubos et al., 2024). Current research focuses on understanding how practitioners aim to increase their productivity and make their work process more efficient by targeting LLMs to automate the generation of software artifacts or receive guidance on how to solve certain problems (Khojah et al., 2024b; Storey and Zagalsky, 2016).
Challenges and limitations of LLMs hinder their effectiveness, such as unhelpful responses, which can lead to frustration among engineers (Weisz et al., 2022). This frustration contributes to techno-stress, affecting their workflow, well-being, and productivity (Tarafdar et al., 2007; Müller et al., 2023). While frustrations have gotten little attention in LLM research for software engineering, we argue that understanding the causes of frustrations when using LLMs for software-related tasks is the first step to minimising them and thus improving the productivity and well-being of practitioners in the SE industry and academia. Moreover, revealing such triggers helps the designers of LLM-powered tools (e.g., AI Chatbots) improve the user experience and evolution of such tools.
This exploratory study aims to empirically investigate the causes of frustrations in software engineers’ interactions with LLMs (and LLM-powered chatbots) and propose strategies for improvement. We focus on the following research questions:
@@ 1. Introduction
Our study presents four main categories that can cause frustrating emotions for the software engineer. We found that the main cause of frustration is when the software engineers receive an unhelpful or incorrect answer, followed by misunderstanding the intention, failing to meet personal preferences, and other limitations of LLMs. We argue that these categories are specific to SE since it is a domain that requires precise, context-aware, and technically accurate responses that can be directly applied, like using LLMs for code generation.
@@
We report that while frustration from unmet expectations can momentarily impact motivation, software engineers typically remain engaged in tasks despite these frustrating interactions with LLMs. This suggests that such interactions are not usually disruptive enough to prevent task completion.
RQ3: How can the user experience be improved to reduce frustration for software engineers?
We provide recommendations for improvements based on software engineers’ expectations and lived experiences. Engineers offer practical, grounded recommendations that reflect user needs and expectations since they are the primary stakeholders directly interacting with LLMs in real-world contexts and can guide chatbot designers in enhancing design and usability.
In addition, we suggest managers provide training and raise awareness among software engineers in order to manage and minimise frustrations.
2. Background and Related Work
@@
This section presents the conceptual framework of this study, as well as, previous work done on the topic and the the research gap addressed.
2.1. Large Language Models in Software Engineering
@@
In SE, recent research has shown that LLMs have the potential to support practitioners in a variety of tasks including, but not limited to implementation, testing, requirements engineering (Dakhel et al., 2023; Wang et al., 2024; Norheim et al., 2024; Khojah et al., 2024b). Moreover, despite the challenges LLMs impose on academia (e.g., bias and hallucination), they provide many opportunities for researchers and educators as assistants with creating study guides and academic writing (Meyer et al., 2023).
Despite this extensive use and the time saved when automating tasks, integrating LLMs into SE practices can lead to user frustration, which can be understood as the emotional state experienced by a person when they are prevented or hindered from obtaining something they have been led to expect (Association, ndb).
@@
Researchers have noted that while LLMs assist software practitioners, they may sometimes generate errors, causing them to spend additional time solving errors or seeking clarification. This can also trigger negative emotions, such as frustration (Weisz et al., 2022).
Our study addresses a gap in the literature by exploring whether the above findings also apply to LLM interactions for SE tasks, with a focus on frustration.
2.2. Emotions Involved When Using Technology
@@
In this study, we considered the definition by the American Psychological Association (Association, nda), Emotion is “a complex reaction pattern, involving experiential, behavioural, and physiological elements, by which an individual attempts to deal with a personally significant matter or event”. We acknowledge the diversity of the conceptual definitions of emotions, we selected this definition for its scope, which integrates experiential, behavioural, and physiological dimensions. The Emotions Wheel, developed by Gloria Willcox (Willcox, 1982), is a psychological tool designed to facilitate identifying and expressing emotions. It is widely used in therapeutic settings and personal development to help individuals articulate their emotional states more precisely. The tool organises emotions into a concentric structure, with six core emotions—happy, sad, angry, scared, strong, and calm—situated at the center of the wheel (See Figure1). As one moves outward from the core, these primary emotions subdivide into more specific and nuanced feelings, allowing for a more granular understanding of emotional experiences.
Figure 1. Willcox’s Emotions Wheel (Neff, 2024)
[…]
@@
In this study, we used the adaptation done by (Neff, 2023) to explore how individuals experience their interaction with LLMs.
While tools like the Emotions Wheel help categorise and articulate emotional experiences, technology-specific contexts introduce unique emotional challenges. These include emotional strain, which refers to adverse reactions and feelings triggered by stressors (Chang et al., 2007), and technostrain and its components, particularly relevant for understanding interactions with LLMs.
@@
Salanova (Salanova et al., 2014) defines Technostrain as a negative psychological experience characterised by i) heightened anxiety and fatigue (affective dimension), ii) scepticism (attitudinal dimension), and iii) a sense of inefficacy (cognitive dimension) associated with technology use. Aligned with this, Muller et al. (Müller et al., 2023) researched techno-frustration, a component of technostrain, which refers to the psychological strain caused by the disorganised or inefficient use of Information and Communications Technology (ICT). Techno-frustration describes the experience of feeling discouraged, uncertain, stressed, confused, and upset as a result of using ICT. Such psychological strain can lead to decreased job satisfaction or an increased risk of burnout (Srivastava et al., 2015). Muller et al.’s work is the closest to our study; however, we focus on a specific interaction, LLM-user interaction, which is still underexplored despite the rapid integration of LLMs in SE.
Regarding research on how technology impacts users’ emotions, studies have focused on the psychological effects of prolonged technology use, recognising that digital tools influence productivity and mental and emotional states (Calvo and Peters, 2014; Mark et al., 2015). Wester et al. (Wester et al., 2024) found that incorrect outputs, which reject the user’s request, can lead to frustration and greatly diminish their perception of the LLM’s usefulness, appropriateness, and relevance.
@@
This reflects an effort to explore the balance between the benefits of technology and the potential strain it places on users, particularly in high-demand environments like SE. Examples of those efforts are in the form of interventions (Montes et al., 2024), (Penzenstadler et al., 2022) or looking for the causes (Wong et al., 2023).
@@
Furthermore, users’ expectations when using technology, particularly LLMs, are crucial to their frustration. Studies have shown that prior experiences with technology can shape users’ expectations, influencing their perceptions of performance and trust (Ferreri and Mayhorn, 2023). When technology does not meet users’ expectations—whether based on past interactions or external portrayals—frustration can escalate, hindering effective task completion and satisfaction  (Srivastava et al., 2015). By examining these responses, we aim to inform LLM design improvements that enhance productivity and reduce technostress.
2.3. Emotions in Software Engineering Tasks
Emotions related to software engineering tasks have been widely researched. Several studies have identified a wide range of emotions, from anger and frustration to joy and satisfaction, as software engineers perform their tasks and communication channels within the development context (Foster and Sayers, 2012; Mäntylä et al., 2014; Graziotin et al., 2014; Marcos et al., 2020; Mäntylä et al., 2016; Russell, 1991; Girardi et al., 2022). Sánchez-Gordón’s (Sánchez-Gordón and Colomo-Palacios, 2019) literature review further analysed the diversity of emotions developers face, identifying 40 discrete emotions, the most frequent being anger, fear, disgust, sadness, joy, love, and happiness. Various situational and contextual factors shape this rich diversity of emotions. Additionally, the impact of affective states has been related to performance and code quality. For example, Graziotin et al. (Graziotin et al., 2013) found that positive emotions, such as happiness, are closely linked to improved performance and productivity. Conversely, negative emotions, such as frustration and unhappiness, can reduce motivation, hinder task completion, and increase the likelihood of turnover (Graziotin et al., 2018). Furthermore, studies have found that developers report higher productivity when in a state of flow and often experience frustration due to being stuck, technical difficulties or unfulfilled information needs (Müller and Fritz, 2015; Girardi et al., 2020).
Moreover, specific triggers of negative emotions have been extensively studied. For example, unhappiness, Graziotin et al. (Graziotin et al., 2017) found that everyday sources of unhappiness are time pressure, bad code quality, repetitive tasks, and inadequate decision-making. Regarding frustration, Ford and Parnin (Ford and Parnin, 2015) identified program comprehension challenges, poor tooling, and fear of failure as common causes.
Despite significant research on emotions associated with traditional software engineering tasks, there is still a gap in investigating the emotional impact of interactions with emerging tools like Large Language Models (LLMs). This gap leaves unexplored how emotions such as frustration evolve in response to LLM use, which has critical implications for improving their design and effectiveness in supporting software engineering tasks.
3. Methodology
Our study aims to understand software engineers’ frustrations and emotions when interacting with LLMs, identifying causes and potential solutions. This study implements an exploratory design to gather initial insights and detect interaction patterns (Stebbins, 2001). We used surveys following Stol and Fitzgerald guidelines (Stol and Fitzgerald, 2020) since they are particularly effective for exploratory studies aiming to generalise findings across a population. We collected qualitative and quantitative data through open and closed questions, covering specific and broader aspects of LLM usage. Figure 2 illustrates our methodology.
Figure 2. The exploratory study process that we followed to design our survey and analyse the responses qualitatively and quantitively.
[…]
[study process]The exploratory study process that we followed to design our survey and analyse the responses qualitatively and quantitively.
3.1. Target Population
We surveyed software engineers in academia and industry to get a broader exploration of how LLMs are used in software engineering tasks. While industry engineers provided information about practical applications such as coding, debugging, and meeting production demands, academics, on the other hand, added LLM usage in research-related tasks such as programming, testing and general research on software processes and quality. Despite differences in their environments, both groups share core practices and challenges, such as dealing with tool limitations and managing cognitive load. Including both populations ensured a comprehensive analysis of how LLMs impact software engineering workflows, making our findings broadly relevant to diverse real-world and research contexts.
3.2. Data Collection
@@
We used a questionnaire as our main data collection method. We designed the survey questions based on previous research about the usage of LLMs in SE and how it impacts their experience and productivity (Khojah et al., 2024b) as well as general limitations of LLMs (Hadi et al., 2023).
@@
After a pilot study with two PhD students (academia) and two software engineers (industry), we added more questions and refined the survey (available in our Zenodo package (Martinez Montes and Khojah, 2025)). The survey started asking for consent to participate and included (i) four demographic questions and queries about participants’ experience and familiarity with LLMs, (ii) two open-ended and one closed-ended question about their usage of LLMs during their work and their expectations from these LLMs, (iii) four open-ended questions and one closed-ended question about the participants’ emotions when certain expectations are not met and about frustrations specifically that the participants experienced when using LLMs, and (iv) five Likert-scales of the level of importance of different LLMs abilities and aspects to minimise frustrating experiences.
@@
The survey was created on Google Forms and distributed across social platforms (LinkedIn and Facebook). We used stratified random sampling to get software engineers from academia and industry, drawing on Baltes and Ralph’s (Baltes and Ralph, 2022) work as a framework for our sampling approach. We sent approximately 20 personal invitations to employees across 10 software organisations of different sizes (Startups and large companies) and domains (e.g., automotive and eLearning). We also invited researchers and academics in SE conferences (RE’24 and FSE’24) to allow our sample of software engineers to be diverse in terms of countries and domains. All data was anonymised, we did not ask for personal data or identification. We followed our university’s ethical regulations and guidelines. We believe that it is important that our sample was diverse in terms of areas and domains, particularly that it included software engineering academics. Software engineering practices often overlap between academia and industry, with some differences in priorities and contexts. This diversity enabled us to capture a broader range of perspectives and demonstrate that academics and practitioners share emotions and challenges related to using LLMs in the field.
3.3. Data Analysis
@@
We used content analysis with an inductive approach following the steps by Erlingsson and Brysiewicz (Erlingsson and Brysiewicz, 2017) to analyse the open questions. Both authors carried out the whole data analysis together systematically. The analysis started by reviewing each answer in detail and discussing them to ensure a shared understanding. This allowed us to identify initial patterns and codes. The codes were then categorised based on their similarities and differences  (Erlingsson and Brysiewicz, 2017), with the categories refined iteratively to ensure accuracy. Themes emerged organically from the data, reflecting the participants’ perspectives and providing meaningful insights. We used the emotion classification and feeling wheel by Willcox (Willcox, 1982) to categorise and identify the range of emotions expressed by participants. For example, one participant’s comment, ”I acknowledge that it might give incorrect answers so it is indifferent for me unless it happens often” was coded as Indiferent’. Enabling a deeper understanding of their emotional responses during interactions with LLMs. For the Likert and close questions, we use descriptive statistics and visualisations.
4. Results
This section presents the results of the visualisation of the closed questions and the content analysis of the open questions. We distinguished between academics and practitioners when their results differed, such as in LLM usage. We combined the results when their patterns were similar, like emotional responses or frustration triggers.
4.1. Respondents Demographics